# Modelling Coincident Particle Production in Ultraperipheral Heavy Ion Collisions

### v1

@@ Abstract
In this paper, we present an analysis of coincident particle production in ultraperipheral heavy ion collisions. In particular, we present the first detailed and differential predictions for coincident $\rho^{0}$ meson production in association with muon pair production, motivated by the recent ATLAS measurement of this process. These are found to describe the data well, including the dependence of the coincidence fraction on the ZDC selection and/or other kinematic constraints on the muons. Differential predictions at the level of various kinematic variables are presented, and the calculation is made publicly available in the SuperChic MC generator. We also present general results for coincident two photon–initiated production, focussing on muon and electron pair production; while the former is rather suppressed, the latter will be ubiquitous near threshold. The impact of coincident production on exclusivity vetos in ultraperipheral measurements is discussed.
@@ 1 Introduction
In this paper, we correct this state of affairs, presenting the first detailed and differential predictions for coincident $\rho^{0}$ meson production in association with lepton pair production. These predictions are implemented and will be made publicly available in the SuperChic Monte Carlo generator. We will focus on the recent ATLAS analysis, and show that indeed these result show a very encouraging level of agreement with theoretical expectations. With this in mind, we will also provide predictions for further differential observables that may be tested in the future.
In addition to this, we will discuss the impact any such coincident production may have on observations of UPC production due to individual photon emission from each ion. Namely, as an exclusivity requirement is imposed in such analyses, a ‘no–coincident’ probability should be accounted for in theoretical predictions (or equivalently corrected for in the data) in a manner that is not currently performed. The impact of this is expected to be small, at the percent level, but may lead to an improvement in the description of not just the overall rate for certain processes but also the ZDC event fractions.
@@ 2.3 Including Coincident Production
i.e. we simply replace the ion dissociation cross section $\gamma A\to A^{*}$ with the exclusive photoproduction cross section $\gamma V\to V$ of a state V. This can be calculated, or evaluated from the relevant direct experimental determinations of the photoproduction process. The factor of 2 here accounts for the fact that the coincident production may be due to photon emission from either ion.
@@
To illustrate the trend with impact parameter, in Fig. 1 (left) we plot the square of (18), i.e. the breakup probabilities for no or multiple neutron emission at $\sqrt{s_{nn}}=5.02$ TeV in PbPb collisions, as a function of the squared ion–ion impact parameter, including the ion–ion survival factor and with/without (shown by the dashed/solid lines) coincident $\rho^{0}$ meson production. The precise method for calculating the $\rho^{0}$ photoprodution cross section is described in the following section, and the rate is multiplied by a factor of 50 in the plot for ease of comparison. Our basic expectation will then be that the relative fraction of the integrand that remains after requiring coincident $\rho^{0}$ production will determine the relative rates at which this occurs between the different dissociation cases. In particular, for the $0n0n$ case, which is most peaked towards higher $b_{\perp}$, we will expect the smallest rate, and for the $XnXn$ case, which is most peaked towards lower $b_{\perp}$, we will expect the largest rate. For other cases, such as $AnAn$, which lie in between, we will likewise expect the coincidence rate to be between these. As we will see, this is indeed observed in the quantitative predictions.
@@
Finally, if instead of coincident photoproduction we are interested in coincident $\gamma\gamma\to V$ production, then this can be accounted for in a rather similar way. In particular, the photoproduction cross section is simply given by integrating the probability (15) over the ion–ion impact parameter, i.e. we have
$\sigma_{A_{1}A_{2}\to A_{1}A_{2}V}=\int{\rm d}^{2}b_{\perp}P_{V}(b_{\perp})\;.$ (23)
@@
$P_{V,\gamma\gamma}(b_{\perp})=\frac{1}{2}\int{\rm d}^{2}b_{\perp}^{\prime}\frac{{\rm d}x_{1}}{x_{1}}\frac{{\rm d}x_{2}}{x_{2}}|\tilde{\mathbf{N}}(x_{1},b_{1\perp})|^{2}|\tilde{\mathbf{N}}(x_{2},b_{2\perp})|^{2}\,\sigma_{\gamma\gamma\to X}({\mathbf{b}}_{1\perp},{\mathbf{b}}_{2,\perp})\;,$ (25)
@@ 3 Results
3.1 Coincident two photon–initiated production: general results
@@
$P_{V,\gamma\gamma}(b_{\perp})=F_{V,\gamma\gamma}(b_{\perp})\cdot\sigma_{V,\gamma\gamma}\;,$ (26)
@@
$\sigma_{V,\gamma\gamma}$ $=\int\frac{{\rm d}x_{1}}{x_{1}}\frac{{\rm d}x_{2}}{x_{2}}\,\sigma_{\gamma\gamma\to X}\;,$ (28)
@@
$\sigma_{V,\gamma\gamma}\sim\frac{8\pi\alpha^{2}}{(M_{ll}^{\rm min})^{2}}\sim\frac{10^{-3}}{(M_{ll}^{\rm min})^{2}}\;.$ (30)
@@
In terms of the level of suppression, we find that the flux component of the coincidence probability is enhanced by roughly two orders of magnitude relative to the photoproduction case, primarily due to the increased $\sim Z^{4}$ scaling of the result in comparison to the $\sim Z^{2}$ scaling of the photoproduction case. That is, at $b_{\perp}=2R_{A}$ we have
@@
$\sigma_{V,\gamma\gamma}\sim\frac{\pi\alpha^{2}}{m_{l}^{2}}\sim 0.02\,(600)\;{\rm GeV}^{-2}\;,$ (32)
for muon (electron) pair production. That is, after accounting for the photon flux suppression as in Fig. 1 (right), the coincidence rate for muon pair production would be at the permille level, while for electron pair production this will become significantly larger than one, and will remain so even beyond threshold. That is, while coincident muon pair production appears to be rather challenging to observe, for coincident electron production we expect multiple pair production to occur for essentially all UPC collisions. This can be readily dealt with in a similar manner to the case of mutual ion dissociation, as discussed in [12], i.e. through a standard process of unitarisation. The key issue is, as discussed in [25], the extent to which such production near the electron mass threshold will be observable, and/or will impact on any exclusivity veto, or rather how far above threshold the production must be for this to be the case. We note that the above estimate for muon pair production is entirely consistent with the calculation of [13], which predicts a contribution of below 1% for multiple muon pair production at threshold, see [25] for more discussion and [14, 17] for other studies.
@@ 3.2 Coincident photoproduction: comparison to ATLAS data
In this section we will present numerical results for the case of coincident photon–initiated muon pair and $\rho^{0}$ meson photoproduction, motivated by the recent ATLAS measurement of this process [8]. This is calculated using the approach outlined in the previous sections, and is implemented in the SuperChic MC generator. This simulates the differential cross section for dimuon production with coincident $\rho^{0}$ photoproduction, for standard ZDC selections, but does not explicitly generate the $\rho^{0}$ meson final–state and its decay products. In particular, the kinematics of the produced $\rho^{0}$, and its decay, can be reasonably assumed to be independent of the dimuon system, and hence can be simulated if necessary (e.g. to evaluate acceptance corrections and so on) but a dedicated MC generator for $\rho^{0}$ photoproduction, such as STARLight [26].
@@
We set the normalization of the $\gamma A\to\rho^{0}A$ subprocess cross section in (15) using the experimental value from the ALICE measurement [27] of $\rho^{0}$ production in UPCs at 5.02 TeV. In particular, as this is expected to depend rather weakly on the photon energy [18, 28] we simply assume a constant cross section. Applying (23) we can then translate this into a UPC cross section; we find a value of 2.3 mb for $\sigma_{\gamma A\to VA}$ gives a cross section in the ALICE case of 560 mb in the central $|y_{\rho^{0}}|<0.2$ bin, in excellent agreement with the experimental value of 537 $\pm$ 4.6 (stat) ${}^{+46.1}_{-42.0}$ (sys) mb. The precise value is chosen so as to match the ATLAS $AnAn$ (no selection) results, up to the issue of the $\rho^{0}$ decay acceptance, which we will discuss below. We note it is broadly consistent with other determinations, see e.g. [28]. Taking this, we then integrate over a photon energy range roughly in line with the ATLAS rapidity selection, namely for $|y_{\rho^{0}}|<2.5$. While this is our baseline, as the coincident cross section is simply linearly dependent on the overall normalization of $\sigma_{\gamma A\to VA}(\omega)$, it is straightforward to vary this when comparing to data, if required.
@@
We first consider the overall fraction of $\mu^{+}\mu^{-}$ UPC events with coincident $\rho^{0}$ production for different ZDC selections, with no additional kinematic requirements applied. The final–state muons are in all cases required to lie within the ATLAS event selection. This is shown in Fig. 2, where a comparison is presented to the ATLAS data [8]. The overall trend for an increasing fraction from the $0n0n$ to the $XnXn$ ZDC selections, as discussed in the previous section, is indeed observed in the predictions, with a sizeable difference seen between these two most extreme cases. In terms of overall normalization of the predictions, some further comment is required. In particular, while the predictions correspond to the $|y_{\rho^{0}}|<2.5$, in the ATLAS analysis results are presented with a somewhat different requirement. Namely, no acceptance correction due to the requirement that the decay pions must lie within the fiducial region ($|\eta_{\pi}|<$ 2.5, $p_{\perp,\pi}>$ 100 MeV) is applied. In [8] it is commented that roughly 76% of $\rho^{0}$ events with $|y_{\rho^{0}}|<2.5$ are found in the ATLAS event selection on the pions from the $\rho^{0}$ decay, when using the STARLight MC generator [26]. In other words, to match the ATLAS selection we should scale our baseline value of $\sigma_{\gamma A\to VA}$ by $1/0.76$, i.e. from 2.3 to 3 mb. This would correspond to an ALICE UPC cross section of about 750 mb, in disagreement with the measurement described above [^1] We note that the 5.02 TeV ALICE data is consistent with the ALICE measurement [29] at 2.76 TeV, i.e. in line with the lower cross section normalization.. The reason for this tension is an open question, but we note that up to this overall normalization the results in Fig. 2 are found to closely match the ATLAS data, within the experimental uncertainties. In particular, this agreement is good at the level of the relative fractions for different ZDC selections, and as we will see below for different kinematic requirements. Given this, and the fact that the overall coincident rate is simply linearly dependent on the $\gamma A\to VA$ cross section (up to any photon energy dependence it may have), the reason for the discrepancy may in part to lie with the overall measured rate for $\rho^{0}$ photoproduction between the two experiments; of possible relevance is the somewhat broader rapidity coverage in the ATLAS case. Connected to this, a more careful treatment of the $\rho^{0}$ decay acceptance, i.e. not effectively treating it as constant with $y_{\rho^{0}}$ as we have with the constant rescaling above, may improve the agreement.
@@
In Fig. 5 we show the dependence on an additional kinematic variable not considered in the ATLAS analysis, namely the dimuon acoplanarity $\alpha=1-\Delta\phi_{\mu\mu}/\pi$. We in particular show the average acoplanarity within the ATLAS selection, for the different ZDC classes. As discussed in [25] and elsewhere, this is particularly sensitive to the peripherality of the interaction, and hence we may expect an impact on the acoplanarity distribution from requiring coincident production. This is indeed observed in the figure, where we can see that by requiring coincident $\rho^{0}$ production, as the average impact parameter is reduced the average photon transverse momentum, and hence dimuon acoplanarity, is increased. The relative increase for the difference ZDC classes is not uniform, in line with the discussion above.
@@
We next consider a different process, namely dilepton production where no coincident process is required. Of particular note here is the fact that in such measurements an exclusivity requirement is made to select the data. Therefore, any coincident particle production of the type discussed above that does not pass the particular experimental veto will be rejected. The overall expectation would therefore be that the theoretical predictions, where no such veto is imposed, will overshoot the data by the coincidence fraction. Interestingly there is in some cases a trend for some overshoot to occur. This is discussed in [12], where it is noted that the SuperChic predictions [19] for the ATLAS data on dimuon production in PbPb collisions [30, 31] tend to overshoot the data by roughly 10%. This is clearly larger than the coincident rates for $\rho^{0}$ production observed in [8], but the precise effect will depend on the difference between the event selection applied there, and the (looser) requirement for a veto, where e.g. only one additional track is required to lie in the fiducial region, as well on other coincident processes beyond $\rho^{0}$ photoproduction that may play a role.

### v2

@@ Abstract
In this paper, we present an analysis of coincident particle production in ultraperipheral heavy ion collisions. In particular, we present the first detailed and differential predictions for coincident $\rho^{0}$ meson production in association with muon pair production, motivated by the recent ATLAS measurement of this process. These are found to describe the data very well, including the dependence of the coincidence fraction on the ZDC selection and/or other kinematic constraints on the muons. Differential predictions at the level of various kinematic variables are presented, and the calculation is made publicly available in the SuperChic MC generator. We also present general results for coincident two–photon initiated production focussing on muon and electron pair production; while the former is rather suppressed, the latter will be ubiquitous at threshold. The impact of coincident production on exclusivity vetos in ultraperipheral measurements is in addition discussed.
@@ 1 Introduction
In this paper, we correct this state of affairs, presenting the first detailed and differential predictions for coincident $\rho^{0}$ meson production in association with lepton pair production. These predictions are implemented and is made publicly available in the SuperChic Monte Carlo generator. We will focus on the recent ATLAS analysis, and find that indeed these results show a very encouraging level of agreement with theoretical expectations. With this in mind, we will also provide predictions for further differential observables that may be tested in the future.
In addition to this, we will discuss the impact any such coincident production may have on observations of UPC production due to individual photon emission from each ion. Namely, as an exclusivity requirement is imposed in such analyses, a ‘no–coincident’ probability should be accounted for in theoretical predictions (or equivalently corrected for in the data) in a manner that is not necessarily currently performed. The impact of this is expected to be small, at the percent level, but may lead to an improvement in the description of not just the overall rate for certain processes but also the ZDC event fractions.
@@ 2.3 Including Coincident Production
i.e. we simply replace the ion dissociation cross section $\gamma A\to A^{*}$ with the exclusive photoproduction cross section $\gamma A\to AV$ of a state V. This can be calculated, or evaluated from the relevant direct experimental determinations of the photoproduction process. The factor of 2 here accounts for the fact that the coincident production may be due to photon emission from either ion.
@@
To illustrate the trend with impact parameter, in Fig. 1 (left) we plot the square of (18), i.e. the breakup probabilities for no or multiple neutron emission at $\sqrt{s_{nn}}=5.02$ TeV in PbPb collisions, as a function of the squared ion–ion impact parameter, including the ion–ion survival factor and with/without (shown by the dashed/solid lines) coincident $\rho^{0}$ meson production. The precise method for calculating the $\rho^{0}$ photoprodution cross section is described in the following section, and the rate is multiplied by a factor of 50 in the plot for ease of comparison. We note that here $Xn$ corresponds to 1 or more neutrons being produced, while $An$ is inclusive with respect to neutron production.
Our basic expectation will then be that the relative fraction of the integrand that remains after requiring coincident $\rho^{0}$ production will determine the relative rates at which this occurs between the different dissociation cases. In particular, for the $0n0n$ case, which is most peaked towards higher $b_{\perp}$, we will expect the smallest rate, and for the $XnXn$ case, which is most peaked towards lower $b_{\perp}$, we will expect the largest rate. For other cases, such as $AnAn$, which lie in between, we will likewise expect the coincidence rate to be between these. As we will see, this is indeed observed in the quantitative predictions.
@@
Finally, if instead of coincident photoproduction we are interested in coincident $\gamma\gamma\to X$ production, then this can be accounted for in a rather similar way. In particular, the photoproduction cross section is simply given by integrating the probability (15) over the ion–ion impact parameter, i.e. we have
$\sigma_{A_{1}A_{2}\to A_{1}A_{2}X}=\int{\rm d}^{2}b_{\perp}P_{X}(b_{\perp})\;.$ (23)
@@
$P_{X,\gamma\gamma}(b_{\perp})=\frac{1}{2}\int{\rm d}^{2}b_{\perp}^{\prime}\frac{{\rm d}x_{1}}{x_{1}}\frac{{\rm d}x_{2}}{x_{2}}|\tilde{\mathbf{N}}(x_{1},b_{1\perp})|^{2}|\tilde{\mathbf{N}}(x_{2},b_{2\perp})|^{2}\,\sigma_{\gamma\gamma\to X}({\mathbf{b}}_{1\perp},{\mathbf{b}}_{2,\perp})\;,$ (25)
@@ 3 Results
3.1 Coincident two–photon initiated production: general results
@@
$P_{X,\gamma\gamma}(b_{\perp})=F_{\gamma\gamma}(b_{\perp})\cdot\sigma_{X,\gamma\gamma}\;,$ (26)
@@
$\sigma_{X,\gamma\gamma}$ $=\int\frac{{\rm d}x_{1}}{x_{1}}\frac{{\rm d}x_{2}}{x_{2}}\,\sigma_{\gamma\gamma\to X}\;,$ (28)
@@
$\sigma_{X,\gamma\gamma}\sim\frac{8\pi\alpha^{2}}{(M_{ll}^{\rm min})^{2}}\sim\frac{10^{-3}}{(M_{ll}^{\rm min})^{2}}\;.$ (30)
@@
In terms of the level of suppression, we find that that the flux component of the coincidence probability is enhanced by roughly two orders of magnitude relative to the photoproduction case, primarily due to the increased $\sim Z^{4}$ scaling of the result in comparison to the $\sim Z^{2}$ scaling of the photoproduction case. That is, at $b_{\perp}=2R_{A}$ we have
@@
$\sigma_{X,\gamma\gamma}\sim\frac{\pi\alpha^{2}}{m_{l}^{2}}\sim 0.02\,(600)\;{\rm GeV}^{-2}\;,$ (32)
for muon (electron) pair production. That is, after accounting for the photon flux suppression as in Fig. 1 (right), the coincidence rate for muon pair production would be at the permille level, while for electron pair production this will become significantly larger than one, and will remain so even beyond threshold. That is, while coincident muon pair production appears to rather challenging to observe, for coincident electron production we expect multiple pair production to occur for essentially all UPC collisions. This can be readily dealt with in a similar manner to the case of mutual ion dissociation, as discussed in [12], i.e. through a standard process of unitarisation. The key issue is, as discussed in [25], the extent to which such production near the electron mass threshold will be observable, and/or will impact on any exclusivity veto, or rather how far above threshold the production must be for this to be the case. We note that the above estimate for muon pair production is entirely consistent with the calculation of [13], which predicts a contribution of below 1% for multiple muon pair production at threshold, see [25] for more discussion and [14, 17] for other studies.
@@ 3.2 Coincident photoproduction: comparison to ATLAS data
In this section we will present numerical results for the case of coincident photon–initiated muon pair and $\rho^{0}$ meson photoproduction, motivated by the recent ATLAS measurement of this process [8]. This is calculated using the approach outlined in the previous sections, and is implemented in the SuperChic MC generator. This simulates the differential cross section for dimuon production with coincident $\rho^{0}$ photoproduction, for standard ZDC selections, but does not explicitly generate the $\rho^{0}$ meson final–state and its decay products. In particular, the kinematics of the produced $\rho^{0}$, and its decay, can be simulated if necessary (e.g. to evaluate acceptance corrections and so on) by a dedicated MC generator for $\rho^{0}$ photoproduction, such as STARLight [26].
@@
We set the normalization of the $\gamma A\to\rho^{0}A$ subprocess cross section in (15) using the experimental value from the ALICE measurement [27] of $\rho^{0}$ production in UPCs at 5.02 TeV. In particular, as this is expected to depend rather weakly on the photon energy [18, 28] we simply assume a constant cross section. Applying (23) we can then translate this into a UPC cross section; we find a value of 2.3 mb for $\sigma_{\gamma A\to VA}$ gives a cross section in the ALICE case of 560 mb in the central $|y_{\rho^{0}}|<0.2$ bin, in excellent agreement with the experimental value of 537 $\pm$ 4.6 (stat) ${}^{+46.1}_{-42.0}$ (sys) mb. This is broadly consistent with other determinations, see e.g. [28]. We note that the precise value is chosen so as to match the ATLAS $AnAn$ (no selection) results, subject to the considerations below, but as noted above, it is entirely consistent with the ALICE data. Taking this, we then integrate over a photon energy range roughly in line with the ATLAS rapidity selection, namely for $|y_{\rho^{0}}|<2.5$. While this is our baseline, as the coincident cross section is simply linearly dependent on the overall normalization of $\sigma_{\gamma A\to VA}(\omega)$, it is straightforward to vary this when comparing to data, if required.
In more detail, in the ATLAS analysis the data for $\rho^{0}\to\pi^{+}\pi^{-}$ in fact also includes a contribution from coherent non–resonant $\pi^{+}\pi^{-}$ photoproduction. This is subtracted in the ALICE data, but not in the ATLAS case, and hence we would expect including coherent $\rho^{0}$ photoproduction alone, with a normalization extracted from the ALICE data, to undershoot the ATLAS measurement of the coincident fractions. A fit to the dipion mass spectrum, including the non–resonant piece, is provided by numerous experiments [29, 30, 31, 32], and integrating over this we find a relative contribution of $\sim 15\%-35\%$ from non–resonant production, depending on the fit/dataset. In general we will expect the precise contribution to depend on the particular beam energies (and types) as well the event selection, and we note that the above hadron–hadron results would in principle require correcting to the $\gamma A$ level for our purposes (even if the mass dependence of the photon flux from the ion will be relatively flat over this limited mass region). Taking the STARLight [26] prediction for the ATLAS event selection, we find a value of $\sim 20\%$ [^1] To be precise to arrive at this the code was suitably modified such that the cross section normalization is tied to the integral over the resonant $\rho^{0}$ Breit–Wigner alone, i.e. so that the non–resonant contribution is not absorbed into this normalization., although the precise number will depend on the underlying fit used, in this case to ZEUS data [29]. Hence, this can only be considered as an estimate, but nonetheless we should increase our effective $\rho^{0}\to\pi^{+}\pi^{-}$ by roughly such a factor.
In addition to the above, while the predictions correspond to $|y_{\rho^{0}}|<2.5$, in the ATLAS analysis results are presented with a somewhat different requirement. Namely, no acceptance correction due to the requirement that the decay pions must lie within the fiducial region ($|\eta_{\pi}|<$ 2.5, $p_{\perp,\pi}>$ 100 MeV) is applied. In [8] it is commented that roughly 76% of $\rho^{0}$ events with $|y_{\rho^{0}}|<2.5$ are found in the ATLAS event selection on the pions from the $\rho^{0}$ decay, when using the STARLight MC generator [26]. In other words, to match the ATLAS selection we should scale our effective value of $\sigma_{\gamma A\to VA}$ by $\sim 0.76$. Combining this decrease with the increase described above due to non–resonant production, we find these to first approximation cancel one another. That is, for the results which follow we can simply apply a prediction for purely resonant $\rho^{0}$ photoproduction, with normalization tied to the ALICE data, and consider that the effect of the pion acceptance and non–resonant production have opposite effects. For future data, a subtraction of the non–resonant contribution would enable a more precise comparison. We also emphasise again that a differing acceptance factor and/or impact of non–resonant $\pi^{+}\pi^{-}$ production can be straightforwardly accounted for by simply scaling all coincident cross sections by the relevant factor.
@@
We first consider the overall fraction of $\mu^{+}\mu^{-}$ UPC events with coincident $\rho^{0}$ production for different ZDC selections, with no additional kinematic requirements applied. All theoretical results are calculated at the higher c.m.s. energy value of 5.36 TeV, where most of the data are taken; the correction to account for the data at lower energy can be safely ignored here. This is shown in Fig. 2, where a comparison is presented to the ATLAS data [8]. The overall trend for an increasing fraction from the $0n0n$ to the $XnXn$ ZDC selections, as discussed in the previous section, is indeed observed in the predictions, with a sizeable difference seen between these two most extreme cases. Moreover, the normalization of the coincident fractions is matched well.
@@
In Fig. 5 we show the dependence on an additional kinematic variable not considered in the ATLAS analysis, namely the dimuon acoplanrity $\alpha=1-\Delta\phi_{\mu\mu}/\pi$. We in particular show the average acoplanrity within the ATLAS selection, for the different ZDC classes. As discussed in [25] and elsewhere, this is particularly sensitive to the peripherality of the interaction, and hence we may expect an impact on the acoplanarity distribution from requiring coincident production. This is indeed observed in the figure, where we can see that by requiring coincident $\rho^{0}$ production, the average impact parameter is reduced and therefore the average photon transverse momentum, and hence dimuon acoplanarity, is increased. The relative increase for the difference ZDC classes is not uniform, in line with the discussion above.
We next consider a different process, namely dilepton production where no coincident process is required. Of particular note here is the fact that in such measurements an exclusivity requirement is made to select the data. Therefore, any coincident particle production of the type discussed above that does not pass the particular experimental veto will be rejected. The overall expectation would therefore be that the theoretical predictions, where no such veto is imposed, will overshoot the data by the coincidence fraction, if it is not appropriately corrected for this effect. Interestingly there is in some cases a trend for some overshoot to occur. This is discussed in [12], where it is noted that the SuperChic predictions [19] for the ATLAS data on dimuon production in PbPb collisions [33, 34] tend to overshoot the data by roughly 10%. This is clearly larger than the coincident rates for $\rho^{0}$ production observed in [8], but the precise effect will depend on the difference between the event selection applied there, and the (looser) requirement for a veto, where e.g. only one additional track is required to lie in the fiducial region, as well on other coincident processes beyond $\rho^{0}$ photoproduction that may play a role.